Tom's Tent Company has total fixed costs of $300,000 per year. The firm's average variable cost is $80 for 10,000 tents. At that level of output, the firm's average total costs equal
$80 $90 $100  $110
Answer #1

Answer : The answer is option D.

Average fixed cost = Total fixed cost / Output level = 300,000 / 10,000 = $30

Average total cost = Average variable cost + Average fixed cost = 80 + 30 = $110

Therefore, option D is correct.